نه 2 5. Let F(x) = - [ « tet-2 + t3 +1 dt, find F(2). tt + 3
Answer #1

x3 fir t-2 = te + t² + 1 tt +3 dt find f(2)=? 1 By using Lebnitzs rule of differentiation of intersals fin) f(x) = g(t) dt t